# Gravitational Wave Spectral Shapes as a probe of Long Lived Right-handed Neutrinos, Leptogenesis and Dark Matter: Global versus Local BLB-L Cosmic Strings

### v1

@@ II.2.3 Type II and III seesaw
The Type-I seesaw is the simplest way to generate light neutrino masses via the seesaw mechanism, but it is not the only one. Neutrino masses can also be generated through the Type-II seesaw Magg:1980ut ; Wetterich:1981bx ; Schechter:1980gr , which introduces an electroweak scalar triplet, and the Type-III seesaw Foot:1988aq ; Ma_1998 ; Ma_2002 , which introduces fermionic triplets. We now turn to the thermal history of these triplet states to examine the conditions under which matter domination could occur. Both the type II and III seesaw mechanisms include a BSM particle which is electroweakly charged. The dominant contribution arises from $s$-channel scattering mediated by the electroweak $Z$ boson:
@@
With $g_{Z}$ being large for many species in the Standard Model ParticleDataGroup:2016lqr ; ParticleDataGroup:2020ssz and summing over the possible diagrams, we cannot have freeze out for temperatures below $10^{18}\rm GeV$ and therefore right-handed neutrino masses below the Planck scale, and as such we can never satisfy condition II (Eq. 10). We conclude we cannot have matter domination from the seesaw heavy state in type II and III for thermal or non-thermal production. This is true for any model with Standard Model interactions; BSM interactions, for instance, mediated by $Z^{\prime}$ in the BLB-L extension, can only exacerbate the problem.
@@ II.3 Condition III: Late decays
where $M_{\rm Pl}$ is the Planck mass. We adopt the Casas-Ibarra parameterisation of the Yukawa matrix Casas_2001 , and it emerges that this is suppressed by the masses of the active neutrinos.
@@
where $m$ and $M$ are the diagonal matrices of light and heavy neutrino masses, respectively, and $\tilde{m_{i}}$ denotes the effective neutrino mass. The matrix $U$ is the standard PMNS matrix, and $R$ is a complex orthogonal matrix. This suppression of the decay rate gives us a fighting chance of sufficiently delaying decays to achieve matter domination. With this parameterisation, the condition for right-handed neutrino matter domination reduces to a simple requirement on the effective neutrino mass.
@@
Using the orthogonality condition $\sum_{j}|R_{ij}|^{2}=1$ one can easily show $\text{Min}(\tilde{m_{i}})=\text{Min}(m_{i})$. Thus, the requirement for right-handed neutrino matter domination reduces to a bound on the lightest active neutrino mass,
@@ II.4 Decay Temperature
The condition on the effective neutrino mass ensures that decays occur after the right-handed neutrino dominates the energy budget. The duration is then written in terms of the number of e-folds of matter domination, $N_{e}$, which depends only on the effective neutrino mass,
$N_{e}=\ln\left(\frac{T_{\rm start}}{T_{\rm end}}\right)\simeq\ln\left(\frac{T_{\rm dom}}{T_{\rm dec}}\right)\simeq\ln\!\left(2.04\times 10^{-5}\,\tilde{m}^{-\tfrac{2}{3}}\right),$ (27)
whereas the onset of matter domination is uniquely determined by the right-handed neutrino mass $M$. The numerics of this is also shown later in Fig. 6.
II.5 During of Early Matter Domination
To determine the duration of the matter–dominated era, we solve the coupled Boltzmann system for the entropy–normalised abundances $Y\equiv n/s$. For a right-handed neutrino undergoing the two–body decay $N\to 2R$, where R denotes relativistic radiation particles, the corresponding equations read
@@ II.5 During of Early Matter Domination
$Y_{N}^{\rm eq}(z)=\frac{45\,g_{N}}{4\pi^{4}g_{*}}\,z^{2}K_{2}(z)\ \quad D(z)=\frac{\Gamma}{zH(z)}\ .$ (30)
With initial conditions $Y_{N}(z_{i})=Y_{N}^{i}$ and $Y_{R}(z_{i})=Y_{R}^{\rm eq}$, we solve the system numerically. As a benchmark, Fig. 5 shows the resulting evolution of the right-handed–neutrino and radiation energy densities for $\tilde{m}=10^{-10}\ \text{eV}$ and $M=10^{9}\ \text{GeV}$, together with the effective equation of state for various effective neutrino mass values illustrating the dynamics of the intermediate matter-dominated period.
@@
Figure 6: Duration of matter domination in terms of number of e-folds. Left: Dependence on the effective neutrino mass. A smaller effective mass delays right-handed neutrino decays, extending the period of matter domination. Right: dependence on both $M_{1}$ and $\tilde{m}$ showing matter domination duration is independent of right-handed neutrino mass.
These results show that the onset of matter domination is uniquely determined by the mass of the right-handed neutrino, whilst the duration is determined by the effective neutrino mass. We also found our analytic bound of effective neutrino mass for right-handed neutrino, Eq. 21, is slightly overestimated also.
@@
$N_{e}(\tilde{m})\;\approx\;0.68\,\ln\!\left(\frac{5.22\times 10^{-8}}{\tilde{m}}-1\right)\ .$ (33)
This result closely matches the analytical estimate of Eq. 27 in the small-$\tilde{m}$ limit, exhibiting the same $\tilde{m}-$dependence but differing by only a factor of $\mathcal{O}(2)$ in the logarithmic prefactor. Combining this relation with the equation for $T_{\rm dom}$, Eq. 31, and the equation above, we obtain a formula for the end of matter domination.
$T_{\rm end}(M,\tilde{m})\;\approx\;4.5\times 10^{-3}\,M\,\left(\frac{5.22\times 10^{-8}}{\tilde{m}}-1\right)^{-0.68}$ (34)
@@ III.1 Gravitational wave spectrum from cosmic strings in standard and non-standard cosmic histories
The spontaneous breaking of a $U(1)$ symmetry, for instance, $U(1)_{\rm B-L}$ whether gauged or global, inevitably leads to the formation of a network of cosmic strings, a kind of topological defect in the early universe Kibble:1976sj ; Vilenkin:2000jqa . As the universe evolves, long strings undergo intercommutation, producing closed string loops that subsequently oscillate and radiate energy Vachaspati:1984gt ; Allen:1991bk ; Gouttenoire:2019kij ; Auclair:2019wcv ; Gouttenoire:2022gwi ; Simakachorn:2022yjy , predominantly via gravitational waves (gauged case) or Goldstone bosons (global case) [^1] A persistent discrepancy remains between the precise predictions obtained from field‑theoretic analyses and those from lattice‑based string simulations; see Refs. Vincent:1997cx ; Matsunami:2019fss ; Hindmarsh:2021mnl ; Blanco-Pillado:2023sap for detailed discussions.. The properties of the long string network are encapsulated by a correlation length $L=\sqrt{\mu/\rho_{\infty}}$, defined through the energy density $\rho_{\infty}$ in long strings, where $\mu$ is the string tension defined as,
@@
$l(\tilde{t})=l_{i}-(\Gamma G\mu+\kappa)(\tilde{t}-t_{i}),$ (36)
where $\Gamma\simeq 50$ Vilenkin:2000jqa ; Vachaspati:1984gt , $\alpha\simeq 0.1$ Blanco-Pillado:2013qja ; Blanco-Pillado:2017oxo , $G$ is the Newton’s constant, and $t_{i}$ is the time of loop formation. The shrinking rate is controlled by two contributions, $\Gamma G\mu$, associated with GW emission, and $\kappa$, associated with Goldstone production. For local strings, $\kappa=0$, and loop decay is dominated by gravitational radiation. In contrast, global strings decay predominantly into Goldstone bosons, with an efficiency $\kappa=\frac{\Gamma_{\rm Gold}}{2\pi}\log(v_{\rm B-L}t)\gg\Gamma G\mu$ , where $\Gamma_{\rm Gold}\simeq 65$ Vilenkin:1986ku ; Gorghetto:2021fsn .
@@
Here, $\rho_{c}$ denotes the critical energy density of the universe, $\mathcal{F}_{\alpha}\simeq 0.1$ is an efficiency factor, and $C_{\rm eff}(t_{i})$ is the loop formation efficiency, computable from the velocity-dependent one-scale model Martins:1996jp ; Martins:2000cs ; Sousa:2013aaa ; Auclair:2019wcv ; Sousa:2020sxs . The integral in Eq. (38) is regulated by two Heaviside functions, $\Theta\left(t_{i}-\frac{l_{*}}{\alpha}\right)\Theta(t_{i}-t_{\rm osc})$, which imposes a high-frequency cut-off at $f_{*}$, beyond which the GW spectrum exhibits a slope $f^{-1/3}$ when summed over a large number of modes. The quantity $t_{\rm osc}=\text{Max}\left[t_{\rm form},\,t_{\rm fric}\right]$ marks the epoch at which the motion of the string network ceases to be friction-dominated, or when loops that could have formed before the formation of the network are eliminated. The parameter $l_{*}$ represents a critical loop length above which GW emission dominates over particle production, as confirmed by high-resolution numerical simulations. Eq. (38) applies to both local and global strings, provided Eq. (37) is used in conjunction with an appropriate choice of $\kappa$.
At high frequencies, under standard cosmological evolution, the GW spectrum from local strings is approximately flat, with an amplitude
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$ParticleDataGroup:2020ssz . Small deviations from flatness may arise due to variations in the number of relativistic degrees of freedom, encapsulated in Gouttenoire:2019kij by
@@
In contrast, the high-frequency part of the GW spectrum from global strings is significantly suppressed and can be approximated as Gouttenoire:2019kij
@@
where the time of maximum emission is
@@
A detailed comparison between the GW signatures of local and global strings can be found in Refs. Gouttenoire:2019kij ; Ghoshal:2023sfa .
Figure 7: The GW spectral shapes for local strings (left panel) and global strings (right panel) are shown with a brief period of early matter domination. The period of matter domination is indicated by a kink and decline in the spectrum, whereas the spectrum is relatively flat for frequencies that correspond to periods of radiation domination. Two features in the each benchmark point chosen, consisting of $f_{\rm dom}$ and $f_{\rm brk}$, determining the start and end of matter dominations respectively (see text for details), benchmarks are chosen in such a manner that in one benchmark (BP1) only one timescale that is the time of end of matter domination, is observable, and in the other one (BP2), both the start and end of matter domination timescales are observable (see text for details).
Following Refs Ghoshal:2023sfa , in scenarios with an early matter-dominated epoch, the otherwise flatter plateau of the GW spectrum takes a spectral turnover at a characteristic break frequency $f_{\rm brk}$. For cusp-dominated loop structures, and when summing over many harmonic modes, the spectrum above this break falls as $\Omega_{\rm GW}(f>f_{\rm brk})\propto f^{-1/3}$. The break frequency for local and global string networks can be estimated as,
@@
However, in scenarios with a brief period of early matter domination, the GW spectrum exhibits a double-step feature that is more prominent for local strings. Interestingly, for local strings, prior to the flat plateau associated with early radiation domination, a characteristic knee feature arises due to loops that formed during the early radiation era but emit GWs during the early matter epoch Ghoshal:2023sfa . The position of this knee can be well approximated as,

### v2

@@ II.2.3 Type II and III seesaw
The Type-I seesaw is the simplest way to generate light neutrino masses via the seesaw mechanism, but it is not the only one. Neutrino masses can also be generated through the Type-II seesaw Magg and Wetterich (1980); Wetterich (1981); Schechter and Valle (1980), which introduces an electroweak scalar triplet, and the Type-III seesaw Foot et al. (1989); Ma (1998); Ma and Roy (2002), which introduces fermionic triplets. We now turn to the thermal history of these triplet states to examine the conditions under which early matter domination could occur. Both the type II and III seesaw mechanisms include a BSM particle which is electroweakly charged. The dominant contribution arises from $s$-channel scattering mediated by the electroweak $Z$ boson:
@@
With $g_{Z}$ being large for many species in the Standard Model Patrignani and others (2016); Zyla and others (2020) and summing over the possible diagrams, we cannot have freeze out for temperatures below $10^{18}\rm GeV$ and therefore right-handed neutrino masses below the Planck scale, and as such we can never satisfy condition II (Eq. 10). We conclude we cannot have early matter domination from the seesaw heavy state in type II and III for thermal or non-thermal production. This is true for any model with Standard Model interactions; BSM interactions, for instance, mediated by $Z^{\prime}$ in the BLB-L extension, can only exacerbate the problem.
@@ II.3 Condition III: Late decays
where $M_{\rm Pl}$ is the Planck mass. We adopt the Casas-Ibarra parameterisation of the Yukawa matrix Casas and Ibarra (2001), and it emerges that this is suppressed by the masses of the active neutrinos.
@@
where $m$ and $M$ are the diagonal matrices of light and heavy neutrino masses, respectively, and $\tilde{m_{i}}$ denotes the effective neutrino mass. The matrix $U$ is the standard PMNS matrix, and $R$ is a complex orthogonal matrix. This suppression of the decay rate gives us a fighting chance of sufficiently delaying decays to achieve early matter domination. With this parameterisation, the condition for right-handed neutrino early matter domination reduces to a simple requirement on the effective neutrino mass.
@@
Using the orthogonality condition $\sum_{j}|R_{ij}|^{2}=1$ one can easily show $\text{Min}(\tilde{m_{i}})=\text{Min}(m_{i})$. Thus, the requirement for right-handed neutrino early matter domination reduces to a bound on the lightest active neutrino mass,
@@ II.4 Decay Temperature
The condition on the effective neutrino mass ensures that decays occur after the right-handed neutrino dominates the energy budget. The duration is then written in terms of the number of e-folds of early matter domination, $N_{e}$, which depends only on the effective neutrino mass,
$N_{e}=\ln\left(\frac{T_{\rm start}}{T_{\rm end}}\right)\simeq\ln\left(\frac{T_{\rm dom}}{T_{\rm dec}}\right)\simeq\ln\!\left(2.04\times 10^{-5}\,\left(\frac{\tilde{m}}{\rm eV}\right)^{-\tfrac{2}{3}}\right),$ (27)
whereas the onset of early matter domination is uniquely determined by the right-handed neutrino mass $M$. The numerics of this is also shown later in Fig. 6.
II.5 Duration of Early Matter Domination
To determine the duration of the early matter–dominated era, we solve the coupled Boltzmann system for the entropy–normalised abundances $Y\equiv n/s$. For a right-handed neutrino undergoing the two–body decay $N\to 2R$, where R denotes relativistic radiation particles, the corresponding equations read
@@ II.5 Duration of Early Matter Domination
$Y_{N}^{\rm eq}(z)=\frac{45\,g_{N}}{4\pi^{4}g_{*}}\,z^{2}K_{2}(z)\ \quad D(z)=\frac{\Gamma}{zH(z)}\ ,$ (30)
where $K_{2}$ is the modified Bessel function of the second kind and with initial conditions $Y_{N}(z_{i})=Y_{N}^{i}$ and $Y_{R}(z_{i})=Y_{R}^{\rm eq}$, we solve the system numerically. As a benchmark, Fig. 5 shows the resulting evolution of the right-handed–neutrino and radiation energy densities for $\tilde{m}=10^{-10}\ \text{eV}$ and $M=10^{9}\ \text{GeV}$, together with the effective equation of state for various effective neutrino mass values illustrating the dynamics of the intermediate matter-dominated period.
@@
Figure 6: Duration of early matter domination in terms of number of e-folds of right-handed neutrino domination, $N_{e}$. Left: Dependence on the effective neutrino mass. A smaller effective mass delays right-handed neutrino decays, extending the period of early matter domination. Right: dependence on both $M$ and $\tilde{m}$ showing early matter domination duration is independent of right-handed neutrino mass.
These results show that the onset of early matter domination is uniquely determined by the mass of the right-handed neutrino, whilst the duration is determined by the effective neutrino mass. We also found our analytic bound of effective neutrino mass for right-handed neutrino, Eq. 21, is slightly overestimated also.
@@
$N_{e}(\tilde{m})\;\approx\;0.68\,\ln\!\left(\frac{5.22\times 10^{-8}\ \rm eV}{\tilde{m}}-1\right)\ .$ (33)
This result closely matches the analytical estimate of Eq. 27 in the small-$\tilde{m}$ limit, exhibiting the same $\tilde{m}-$dependence but differing by only a factor of $\mathcal{O}(2)$ in the logarithmic prefactor. Combining this relation with the equation for $T_{\rm dom}$, Eq. 31, and the equation above, we obtain a formula for the end of early matter domination  [^1] Unless otherwise specified, to avoid cumbersome notation the reader should assume any unspecified units are in GeV..
$T_{\rm end}(M,\tilde{m})\;\approx\;4.5\times 10^{-3}\,M\,\left(\frac{5.22\times 10^{-8}\ \rm eV}{\tilde{m}}-1\right)^{-0.68}$ (34)
@@ III.1 Gravitational wave spectrum from cosmic strings in standard and non-standard cosmic histories
The spontaneous breaking of a $U(1)$ symmetry, for instance, $U(1)_{\rm B-L}$ whether gauged or global, inevitably leads to the formation of a network of cosmic strings, a kind of topological defect in the early universe Kibble (1976); Vilenkin and Shellard (2000). As the universe evolves, long strings undergo intercommutation, producing closed string loops that subsequently oscillate and radiate energy Vachaspati and Vilenkin (1985); Allen and Shellard (1992); Gouttenoire et al. (2020a); Auclair and others (2020); Gouttenoire (2022); Simakachorn (2022), predominantly via gravitational waves (gauged case) or Goldstone bosons (global case) [^2] A persistent discrepancy remains between the precise predictions obtained from field‑theoretic analyses and those from lattice‑based string simulations; see Refs. Vincent et al. (1998); Matsunami et al. (2019); Hindmarsh et al. (2021); Blanco-Pillado et al. (2023) for detailed discussions.. The properties of the long string network are encapsulated by a correlation length $L=\sqrt{\mu/\rho_{\infty}}$, defined through the energy density $\rho_{\infty}$ in long strings, where $\mu$ is the string tension related to the symmetry breaking scale as Bogomolny (1976),
@@
$l(t)=l_{i}-(\Gamma G\mu+\kappa)(t-t_{i}),$ (36)
where $\Gamma\simeq 50$ Vilenkin and Shellard (2000); Vachaspati and Vilenkin (1985), $\alpha\simeq 0.1$ Blanco-Pillado et al. (2014); Blanco-Pillado and Olum (2017), $G$ is the Newton’s constant, and $t_{i}$ is the time of loop formation. The shrinking rate is controlled by two contributions, $\Gamma G\mu$, associated with GW emission, and $\kappa$, associated with Goldstone production. For local strings, $\kappa=0$, and loop decay is dominated by gravitational radiation. In contrast, global strings decay predominantly into Goldstone bosons, with an efficiency $\kappa=\frac{\Gamma_{\rm Gold}}{2\pi}\log(v_{\rm B-L}t)\gg\Gamma G\mu$ , where $\Gamma_{\rm Gold}\simeq 65$ Vilenkin and Vachaspati (1987); Gorghetto et al. (2021).
@@
Here, $\rho_{c}$ denotes the critical energy density of the universe, $\mathcal{F}_{\alpha}\simeq 0.1$ is an efficiency factor, and $C_{\rm eff}(t_{i})$ is the loop formation efficiency, computable from the velocity-dependent one-scale model Martins and Shellard (1996a, 2002); Sousa and Avelino (2013); Auclair and others (2020); Sousa et al. (2020). The quantity $\Gamma_{k}=\Gamma k^{-\delta}/\zeta(\delta)$ parametrizes the fraction of GW power emitted into the $k$-th harmonic mode, where $\delta=4/3(5/3)$ for loops containing cusps (kinks)Damour and Vilenkin (2001). The integral in Eq. (38) is regulated by two Heaviside functions, $\Theta\left(t_{i}-\frac{l_{*}}{\alpha}\right)\Theta(t_{i}-t_{\rm osc})$, which imposes a high-frequency cut-off at $f_{*}$, beyond which the GW spectrum exhibits a slope $f^{-1/3}$ when summed over a large number of modes. The quantity $t_{\rm osc}=\text{Max}\left[t_{\rm form},\,t_{\rm fric}\right]$ marks the epoch at which the motion of the string network ceases to be friction-dominated, or when loops that could have formed before the formation of the network are eliminated. The string network forms at the temperature of the $U(1)$ breaking phase transition,
$T_{\rm form}\simeq v_{\rm B-L},$ (39)
which corresponds, during radiation domination, to the formation time
$t_{\rm form}\simeq\left(\frac{90}{\pi^{2}g_{*}(T_{\rm form})}\right)^{1/2}\frac{M_{\rm Pl}}{2T_{\rm form}}.$ (40)
At early times, the motion of strings is damped by interactions with the surrounding plasma. The temperature at which friction becomes inefficient can be estimated as Vilenkin (1991); Martins and Shellard (1996b, a)
$T_{\rm fric}=\left(\frac{\pi^{2}g_{*}(T_{\rm fric})}{90}\right)^{1/2}\frac{2\mu}{\beta M_{\rm Pl}},$ (41)
where $\beta\sim\mathcal{O}(1)$ parametrizes the strength of particle–string interactions. The corresponding friction cutoff time follows from the standard time–temperature relation during radiation domination. Plugging Eq. (41) into Eq. (40), we obtain the time of friction cut-off as
$t_{\rm fric}=\frac{2\pi v_{\rm B-L}}{\beta}\log(v_{\rm B-L}\tilde{t}_{M})\left(\frac{g_{*}(T_{\rm fric})}{g_{*}(T_{\rm form})}\right)^{1/2}.$ (42)
The parameter $l_{*}$ represents a critical loop length above which GW emission dominates over particle production, as confirmed by high-resolution numerical simulations. For local strings, $l_{\star}$ corresponds to the onset of efficient massive particle emission from the string core and can be estimated as
$l_{\star}\sim\delta_{w}(\Gamma G\mu)^{-\gamma},$ (43)
where $\delta_{w}\sim(\sqrt{\lambda}v_{\rm B-L})^{-1}$ is the string width, with a scalar self coupling $\lambda$, and $\gamma=2(\gamma=1)$ for loops containing cusps (kinks) Matsunami et al. (2019); Auclair et al. (2020); Baeza-Ballesteros et al. (2025). For global strings, loop decay is instead dominated by Goldstone boson emission, and no sharp microscopic cutoff exists Baeza-Ballesteros et al. (2024). In this case, $l_{\star}$ should be understood as an effective scale obtained by equating GW and Goldstone emission rates, giving
$l_{*}\sim\kappa(\Gamma G\mu)^{-1},$ (44)
which suppresses the GW contribution from small loops. Eq. (38) applies to both local and global strings, provided Eq. (37) is used in conjunction with an appropriate choice of $\kappa$.
At high frequencies, under standard cosmological evolution, the GW spectrum from local strings is approximately flat  [^3] Ref.Schmitz and Schröder (2024) points out that cutoff or oscillatory features in the GW spectrum can arise in scenarios where GW emission from string loops becomes effective only after a delayed onset and is dominated by the fundamental oscillation mode. In the framework adopted here, GW emission is continuous once loops form and the spectrum is obtained by summing over harmonic modes, so such features are not generically expected., with an amplitude
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$Zyla and others (2020). Small deviations from flatness may arise due to variations in the number of relativistic degrees of freedom, encapsulated in Gouttenoire et al. (2020a) by
@@
In contrast, the high-frequency part of the GW spectrum from global strings is significantly suppressed and can be approximated as Gouttenoire et al. (2020a)
@@
where $\tilde{t}_{M}$ is the characteristic emission time that dominantly contributes to the present–day GW spectrum at a fixed observed frequency $f$. It arises from the interplay between loop shrinkage and redshifting, rather than from a maximum of the instantaneous GW power. For local strings, the GW power is constant in time, and this dominant time dependence cancels in high frequency regime. In contrast, for global strings GW emission is further suppressed by efficient Goldstone radiation, making early emission times particularly relevant. In practice, $\tilde{t}_{M}$ is obtained by maximizing the integrand in Eq. (38) with respect to $\tilde{t}$, leading to
@@
A detailed comparison between the GW signatures of local and global strings can be found in Refs. Gouttenoire et al. (2020a); Ghoshal et al. (2023a).
Figure 7: The GW spectral shapes for local strings (left panel) and global strings (right panel) are shown with a brief period of early matter domination. The period of early matter domination is indicated by a kink and decline in the spectrum, whereas the spectrum is relatively flat for frequencies that correspond to periods of radiation domination. Two features in the each benchmark point chosen, consisting of $f_{\rm dom}$ and $f_{\rm brk}$, determining the start and end of early matter dominations respectively (see Eq.(49),(50),(54),(56)), benchmarks are chosen in such a manner that in one benchmark (BP1) only one timescale that is the time of end of early matter domination, is observable, and in the other one (BP2), both the start and end of early matter domination timescales are observable (see Table 2).
Following Refs. Ghoshal et al. (2023a), in scenarios with an early matter-dominated epoch, the otherwise flatter plateau of the GW spectrum takes a spectral turnover at a characteristic break frequency $f_{\rm brk}$. For cusp-dominated loop structures, and when summing over many harmonic modes, the spectrum above this break falls as $\Omega_{\rm GW}(f>f_{\rm brk})\propto f^{-1/3}$. The break frequency for local and global string networks can be estimated as,
@@
However, in scenarios with a brief period of early matter domination, the GW spectrum exhibits a double-step feature that is more prominent for local strings. Interestingly, for local strings, prior to the flat plateau associated with early radiation domination, a characteristic knee feature arises due to loops that formed during the early radiation era but emit GWs during the early matter epoch Ghoshal et al. (2023a). The position of this knee can be well approximated as,